Dear Tutor this is the problem:
Find three consecutive integers whose sum is 37 less than 5 times the third integer.
I need this at 8:00 tonight please work fast.

Answer by calbreez(9)

x + (x+1) + (x+2) = 5(x+2)-37
3x+3=5x+10-37
3x+3=5x-27
30=2x
x=15
So the integers are 15, 16 and 17